Please help me solve this equation 3/4=5/8 x
Answer by Theo(13342)

I think your formula is:

(3/4) = (5/8) * x

If so, then divide both sides of that equation by (5/8) to get:

(3/4) / (5/8) = x

(3/4) / (5/8) is the same as (3/4) * (8/5) is the same as (3*8)/(4*5).

(3*8)/(4*5) reduces to (3*2)/5 which equals 6/5

(6/5) should be equal to x which would be your answer.

substitute in the original equation to see if the equation is true.

(3/4) = (5/8) * x becomes:

(3/4) = (5/8) * (6/5)

simplify that to get:

(3/4) = (5*6)/(8*5) = (30/40) = (3/4)

since the equation is true, your value of x = (6/5) must be good.